OpenStudy (anonymous):

Solve for x x/10 = 2 1/5 A. 20 1/5 B. 11/50 C. 22 D. 12 1/5

OpenStudy (texaschic101):

x/10 = 2 1/5 -- turn to improper fraction x/10 = 11/5 -- multiply both sides by 10 x = 11/5 * 10 x = 110/5 which reduces to 22 (C)